How to solve this question'
Find the equation of a line perpendicular to y  4x 3, passing through the point (2, 7)
1
Expert's answer
2020-06-14T18:15:29-0400
1.Let the line pass through the point T1 (x1; y1)
 and perpendicular to the line y = kx + b,
 then it can be represented by the equation: y-y1 = -1 / k (x-x1)

from the conditions x1 = 2, y1 = 7,  b=3 



then get y-7 = -1 / 4 (x - 2) or y = o.25x + 7 - 0.5